Acton Company has two products: A and B. The annual production and sales of Product A is 800 units and of Product B is 500 units. The company has traditionally used direct labor-hours as the basis for applying all manufacturing overhead to products. Product A requires 0.3 direct labor-hours per unit and Product B requires 0.2 direct labor-hours per unit. The total estimated overhead for next period is $92,023.

The company is considering switching to an activity-based costing system for the purpose of computing unit product costs for external reports. The new activity-based costing system would have three overhead activity cost pools—Activity 1, Activity 2, and General Factory—with estimated overhead costs and expected activity as follows:

Activity

Estimated Overhead

Costs Expected Activity

Cost Pool

Costs

Product A

Product B

Total

Activity 1

$14,487

500

600

1,100

Activity 2

$64,800

2,500

500

3,000

General Factory

$12,736

240

100

340

   Total

$92,023

(Note: The General Factory activity cost pool's costs are allocated on the basis of direct labor-hours.)

The predetermined overhead rate under the traditional costing system is closest to:

Select one:

a. $37.46.

b. $21.60.

c. $13.17.

d. $270.66.


The overhead cost per unit of Product B under the traditional costing system is closest to:

Select one:

a. $54.13.

b. $7.49.

c. $4.32.

d. $2.63.
Answer #1

Answers :

D) $270.66

A) $54.13

Traditional Costing

Predetermined overhead Rate Under Traditional Costing System:

Rate = Estimated Total cost / Estimated Total Labor Hours

Predetermined Overhead Rate = 92,023 /340, = $270.66

Overhead cost Per Unit for Product B

Unit Cost = applied overhead / Direct labor hours

Units Cost = (100*270.66) /500 = $54.13